My Mum was 27 when I was born.
8 years ago she was twice as old as I shall be in 5 years time.
How old am I?

Answer by rapaljer(4671)

Let x = my age now
x+27 = Mum's age now
x+5= my age in 5 years

Mum's age now - 8 years = Two times my age in 5 years
x+27 - 8 = 2(x+5)
x+19=2x+10

Subtract x from each side:
x+19-x=2x-x+10
19=x+10

Subtract 10 from each side:
9=x

My age is 9 years NOW, and Mum's age is 9+27=36 years NOW.

Check:
Eight years ago Mum was 28. My age in 5 years will be 14. Indeed, her age eight years ago was twice my age in 5 years.

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus